40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 26 22:24:28 2014, Jenne, Update, LSC, DARM loop improved, other work 8x
    Reply  Mon Dec 1 17:43:22 2014, Jenne, Update, LSC, Reset Yarm trans normalization 
Message ID: 10743     Entry time: Mon Dec 1 17:43:22 2014     In reply to: 10737
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Reset Yarm trans normalization 

After Koji and I reset the transmission normalizations last Friday, he did some alignment work that increased the Yarm power.  So, I had set the transmission normalization when we weren't really at full Yarm power.  Today I reset the normalization so that instead of ~1.2, the Y transmission PDs read ~1.0.

ELOG V3.1.3-